# Multi-unit soft sensing permits few-shot learning in virtual flow metering

### v2

@@ 7 Concluding remarks
Fig. 5 suggests that the convergence rate of $\mathcal{O}(1/\sqrt{M})$, where $M$ is the number of units, applies well to the VFM problem. The flatness of this convergence rate for large values of $M$ serves as motivation for our proposed few-shot learning approach, where a pretrained model with fixed parameters acts as a base model. The result is a data-efficient learning algorithm, where only the low-dimensional context parameters are calibrated to data from a new unit.
@@
From Fig. 7, one can see that the dataset information gain varies smoothly with dataset size. This indicates that the regularization which is introduced by the prior distribution over the context variable reduces the complexity of the learned context parameter space. The smoothness of the parameter space as well as its low dimension suggests that the learned context parameters may be meaningful to the model, even if they do not permit a physical interpretation.
@@
Our main result is shown in Fig. 6. From this figure one can see that our proposed calibration procedure only requires a handful of training data points for the resulting model to achieve high performance on a previously unseen well – the median error across new units drops below 5% after calibration to only three data points, which can be considered as a high performance for a VFM [42, 41].
@@
It should be noted that the high performance of the fine-tuned VFM relies on the existence of a good pretrained model. One could argue that we are not in a data-scarce setting, since successful pretraining requires a base dataset with data from a quite large number of well-instrumented wells. However, we emphasize that this pretraining only has to happen once. As long as one has access to a fixed set of wells that give rise to high-quality training data, the few-shot learning methodology can be used to transfer this information to any other well of the same type.
@@
As illustrated by the experiment of Section 6.3, adaptation of the pretrained model to previously unseen wells requires little or no extra manual work. In contrast, the effort required to apply a first principles VFM in the same setting (in which the model must be adapted to 20 new wells) would be prohibitively large. Thus, our proposed fine-tuning strategy may enable the use of VFMs in low-data settings that were previously deemed impractical, e.g. due to insufficient budgets for VFM deployment and maintenance.
@@ Acknowledgments
Appendix A Implementation details
This section describes in further detail how the model in Section 4 and learning algorithm in Section 5 were implemented in the case study of Section 6.
@@ A.1 Model
The multi-unit soft sensor was modeled using a deep neural network, implemented in PyTorch [43]. The network architecture is illustrated in Fig. 8. It consists of a feed-forward neural network of width 400 and depth 4, with ReLU activation functions. Its input is the concatenation of the input data point $x_{ij}$ and a learnable, unit-specific context vector $c_{i}$ of dimension 10. The context vectors of the units can be thought of as learned embedding vectors. Given $x_{ij}$, the neural network returns an estimate $\hat{y}_{ij}=f(x_{ij};c_{i},\theta)$ of the mean of $y_{ij}$. Together with a precision parameter acquired by passing the learnable unit-specific parameter $t_{i}$ through the softmax function $g$, this estimate is used to form the conditional distribution
@@ A.2 Pretraining
During pretraining, all model parameters $(\bm{c},\bm{t},\theta)$ were learned jointly through the mini-batch gradient ascent procedure described in Section 5.1, using the Adam optimizer [44] with a batch size of 2048. For both the convergence rate experiment described in Section 6.2 and the few-shot learning experiment described in Section 6.3, training was run for 20 000 epochs with a learning rate of $1\cdot 10^{-4}$. In both cases, the model with the lowest validation set MSE was returned. The training was conducted on a desktop computer with a single GPU.
@@ A.3 Calibration
Calibration of context parameters $c_{k}$ to previously unseen units $k>M$ was done using the procedure described in Section 5.2. For each new data point, the context parameter $c_{k}$ was reinitialized and training was done from scratch. This training consisted of 100 epochs of stochastic gradient ascent with a learning rate of $1\cdot 10^{-4}$. Due to the low data availability, no model selection procedures were performed in this setting.
Appendix B Model validation in few-shot learning
The dataset described in Section 6.1 consists of data from wells with significant differences in instrumentation and practice of operation. As can be seen in Fig. 4, this results in significant differences between wells in data frequency as well as dataset duration. To increase the comparability of results across wells, we devised a procedure inspired by how new well tests typically arrive in a petroleum production setting. The result was the following procedure, which was used for each of the unseen wells $k>M$:
@@ Appendix B Model validation in few-shot learning
In this procedure, the division into training, validation and test sets shown in Fig. 4 was ignored, and all data from each given well was used when running the procedure described above.
@@
Appendix C Additional results
Fig. 9 shows each of the 20 individual pretraining experiments. The variance observed between individual experiments supports the large number of experiment repetitions performed when generating Fig. 5.
Figure 9: MTL performance for increasing number of $M$. Median across runs is shown in black, individual runs are shown in gray.

### v3

@@ 7 Concluding remarks
Figure 7 suggests that the convergence rate of $\mathcal{O}(1/\sqrt{M})$, where $M$ is the number of units, applies well to the VFM problem. The flatness of this convergence rate for large values of $M$ serves as motivation for our proposed few-shot learning approach, where a pretrained MTL model with fixed parameters acts as a base model. The result is a data-efficient learning algorithm, where only the low-dimensional context parameters are calibrated to data from a new unit.
@@
From Figure 9, one can see that the dataset information gain varies smoothly with dataset size. This indicates that the regularization which is introduced by the prior distribution over the context variable reduces the complexity of the learned context parameter space. The smoothness of the parameter space as well as its low dimension suggests that the learned context parameters may be meaningful to the model, even if they do not permit a physical interpretation. Furthermore, the results shown in Figure 10 indicate that even though a few dimensions accounts for a large part of the modeled inter-unit variation, all $K=10$ context parameter dimensions are exploited to provide a more refined representation of the variation between wells; arguably then, the effective dimension of the context parameter (in the context of finetuning) is $k=K=10$. Given the regularizing role played by the context parameter, this is perhaps not too surprising.
@@
Our main result is shown in Figure 8. From this figure one can see that our proposed calibration procedure only requires a handful of training data points for the resulting model to achieve high performance on a previously unseen well – the median error across new units drops below 5% after calibration to only three data points, which can be considered as a high performance for a VFM [52, 45]. There is still a gap to the performance of about 2,5% error achieved in the pretraining setting, as indicated by the dashed line in Figure 8. We hypothesize that this is due to the non-stationarity of the petroleum production process [24]; over time, a petroleum well depletes and production equipment is subject to wear, meaning the number of up-to-date data points is limited by data observation frequency. Since this frequency is lower in the few-shot setting considered here (where a new data point is observed every week) than in the multi-unit pretraining setting (where about half of the wells have MPFMs and may get access to new data points multiple times a week), a performance gap is to be expected.
@@
It should be noted that the high performance of the fine-tuned VFM relies on the existence of a good pretrained MTL model. One could argue that we are then not in a data-scarce setting, since successful pretraining requires a base dataset with data from a quite large number of well-instrumented wells. However, we emphasize that this pretraining only has to happen once. As long as one has access to a fixed set of wells that give rise to high-quality training data, the few-shot transfer learning methodology can be used to transfer this information to any other well of the same type.
It should also be noted that since the learning problem is highly heterogeneous, the reported prediction errors are aggregated over a large range of different operating conditions and flow regimes. We emphasize that the reported results are an average, and that model performance may co-vary with all of the conditions which results are aggregated over. Even though our results indicate strong model performance on average, it would be interesting to investigate in further detail under which conditions the proposed methodology is the most useful, and under which conditions it struggles. It would be particularly interesting to investigate how model performance is correlated with flow regime. Such an investigation is, however, not straight-forward to conduct, since flow regime classification constitutes its own soft sensing problem. Addressing this soft sensing problem would be an interesting topic for future research, and a solution to this problem could be used both to improve assessment of VFM performance, and to improve the performance of the VFM itself (by providing additional contextual information, e.g. in the form of an additional, data-dependent context variable).
@@
As illustrated by the few-shot learning experiment, adaptation of the pretrained model to previously unseen wells requires little or no extra manual work. In contrast, the effort required to apply a first principles VFM in the same setting (in which the model must be adapted to 20 new wells) would be prohibitively large. Thus, our proposed fine-tuning strategy may enable the use of VFMs in low-data settings that were previously deemed impractical, e.g. due to insufficient budgets for VFM deployment and maintenance.
@@ 7.1 Acknowledgments
7.2 Conflict of interest
Authors state no conflict of interest.
Appendix A Appendix
This section describes in further detail how the model and learning algorithm were implemented and evaluated in the case study.
@@ A.1 Model
The multi-unit soft sensor was modeled using a deep neural network, implemented in PyTorch [53]. The network architecture is illustrated in Figure 11. It consists of a feed-forward neural network of width 400 and depth 4, which has been shown to be sufficiently large in similar studies on neural network-based VFMs (see e.g. [23], [24]). It uses ReLU activation functions, and its input is the concatenation of the input data point $x_{ij}$ and a learnable, unit-specific context vector $c_{i}$ of dimension 10. The context vectors of the units can be thought of as learned embedding vectors. Given $x_{ij}$, the neural network returns an estimate $\hat{y}_{ij}=f(x_{ij};c_{i},\theta)$ of the mean of $y_{ij}$. Together with a precision parameter acquired by passing the learnable unit-specific parameter $t_{i}$ through the softmax function $g$, this estimate is used to form the conditional distribution
@@ A.2 Pretraining
During pretraining, all model parameters $(\bm{c},\bm{t},\theta)$ were learned jointly through the mini-batch gradient ascent procedure described in Algorithm 1, using the Adam optimizer [54] with a batch size of 2048. For both the convergence rate experiment described and the few-shot learning experiment, training was run for 20 000 epochs with a learning rate of $1\cdot 10^{-4}$. In both cases, the model with the lowest validation set MSE was returned. The training was conducted on a desktop computer with a single GPU.
@@ A.3 Calibration
Context parameters $c_{k}$ were calibrated using data from previously unseen units $k>M$. For each new data point, the context parameter $c_{k}$ was reinitialized and training was done from scratch. This training consisted of 100 epochs of stochastic gradient ascent with a learning rate of $1\cdot 10^{-4}$. Due to the low data availability, no model selection procedures were performed in this setting.
A.4 Model validation in few-shot learning
The dataset shown in Figure 6 consists of data from wells with significant differences in instrumentation and practice of operation. As can be seen in Figure 6, this results in significant differences between wells in data frequency as well as dataset duration. To increase the comparability of results across wells, we devised a procedure inspired by how new well tests typically arrive in a petroleum production setting. The result was the following procedure, which was used for each of the unseen wells $k>M$:
@@ A.4 Model validation in few-shot learning
In this procedure, the division into training, validation and test sets shown in Figure 6 was ignored, and all data from each given well was used when running the procedure described above.
@@
A.5 Additional results
Figure 12 shows each of the 20 individual pretraining experiments. The variance observed between individual experiments supports the large number of experiment repetitions performed when generating Figure 7.
Figure 12: MTL performance for increasing number of $M$. The x-axis shows the number of available data points for the wells on which few-shot transfer learning is performed. Median across runs is shown in black, individual runs are shown in gray.